2. A dental clinic has 2400 procedures performed in 2014 and total input cost of $ 750,000. The clinic performs 2250 procedures in 2018 and total input cost of $ 695,000. The PPI (producer price index) for 2014 is 125 (base 100 in 2000) and PPI for 2018 is 139.

What is the % change in MFP adjusted for PPI.
Answer #1

MFP stands for multi-factor productivity.

Multifactor productivity is a measure of the changes in factors of input other than capital and labour, for example : change in management or a change in the same management's practices and therefore its effect on the productivity of the same labour and capital. Similarly, imperfect competition, general knowledge and its effect, measurement errors, also form some other factors which are not categorized under capital and labour.

Therefore, consider the below table which is a summary of the data given in the question and the information available from it.

A B C
Year Total cost of procedures Total number of procedures Cost per procedure(A/B)
2014 750000 2400 312.5
2018 695000 2250 308.9

Considering that the PPI base in 2000 was 100:

2014: To adjust the % change in PPI as 25%, we reduce the cost per procedure by 25% , i.e. 312.5/1.25 =250

2018: Hence, the real price free of any MPF variation should be 250 increased by 39%, i.e. 250*1.39 =347.5

However, the actual cost available from the data is 308.9

This shows that the other factors of production which are not a part of labour or of capital reduced the cost of each procedure.

In the case of a dental clinic, one may consider technology improvement as a significant factor.

Hence, We compare 347.5 and 308.9 to find out the % change in MPF. 308.9/347.5 = 0.8889.

i.e. (1-0.8889)*100 =11.11% i.e. change in cost of performing procedure not caused by labour and capital.